you were in a field counting chickens and rabbits and counted 7 heads and 20 legs. How many chickens were there? How many rabbits were there? Explain how you solved this.
Answer #1

Let no. of chickens = 20 no. of rabbits = y Total heads of chickens and rabbits = x+y x+y = 7 - ul. Total legs of chickens an​​​​​